# Study on physical properties and characteristics of an anisotropic compact star model using Karmarkar Condition in $\mathcal{F}(\mathcal{Q})$ gravity

### v1

@@ II.1 Field equations of $\mathcal{F}(\mathcal{Q})$ gravity
Using equations Eq. (25) and Eq. (26) with $\delta_{1}$ and $\delta_{2}$ as integration constants, the field equations for $\mathcal{F}(\mathcal{Q})$ gravity can be stated as follows:
@@
$p_{r}=\frac{1}{16\pi r^{2}}\Bigg{[}-2\delta_{1}+2e^{-\sigma}\delta_{1}(r\Upsilon^{\prime}+1)+r^{2}\delta_{2}\Bigg{]},$ (28)
$p_{t}=\frac{e^{-\sigma}}{32\pi r}\Bigg{[}2e^{\sigma}r\delta_{2}+\delta_{1}(2+r\Upsilon^{\prime})(\Upsilon^{\prime}-\sigma^{\prime})+2r\delta_{1}\Upsilon^{\prime\prime}\Bigg{]}.$ (29)
In Eqs. (2729) prime ()($\prime$)( ′ ) denotes the differentiation with respect to radial coordinate r and $\rho$, $p_{r}$, $p_{t}$, $\sigma$ and $\Upsilon$ are unknowns which we are going to solve to get our desired result. For static and spherically symmetric fluid configurations, the general expression of the anisotropic factor, denoted as $\Delta=(p_{t}-p_{r})$, can be stated as follows:
@@ II.2 Karmarkar Condition
The spherically symmetric line element (18) can always be embedded in six-dimensional flat spacetime, signifying its classification as belonging to embedding class two in general. On the contrary, the spherical line element can be included in five-dimensional flat spacetime, provided it fulfils the Karmarkar criterion  karmarkar1948gravitational . Then It indicates the spacetime of embedding class one. It is a necessary and sufficient condition for spherically symmetric spacetime to belong to class one. The Karmarkar condition is expressed through the curvature components as follows:
@@ III.1 Metric elements
Anisotropic compact stars require finite and nonsingular metric elements to produce a physically plausible model. In our model, the value of $e^{\sigma}$ at $r=0$ equals 1, and the value of $e^{\Upsilon}$ at r=0 equals 0.4583. we have a positive value of $e^{\sigma}$ and $e^{\Upsilon}$ at r=0, indicating that the metric elements are stable at the centre of the compact star. The Exact behaviour of the metric elements has been shown in Fig. 1. The graph of metric elements $e^{-\sigma}$ and $e^{\Upsilon}$ are finite and nonsingular throughout the radius of the compact star. The graph of $e^{-\sigma}$ is monotonically decreasing towards the surface and the graph of $e^{\Upsilon}$ monotonically increases towards the surface. Both the graph coincides at the boundary of the compact star from these characteristics of metric elements we can conclude that metric elements are competent to produce the model for the anisotropic compact star.
@@ III.2 Seamless Matching of Interior and Exterior spacetime
The boundary requirement for solving the Einstein-field equations in $\mathcal{F}(\mathcal{Q})$ gravity is presented in this section. To accomplish this, we compare the solution on the exterior and interior spacetimes. There are numerous indications in more recent works, such as Wang:2021zaz , that the feasible solution in $\mathcal{F}(\mathcal{Q})$ gravity is solvable using Schwarzchild (Anti-) di sitter solution. The solution is provided as follows:
@@
Taking $r=R$, the Schwarzchild (Anti-) di sitter metric has been solved at the surface using the requirement that Darmois-Israel derived in their research  Israel:1966rt
@@
The analysis from Eq. (46) and Eq. (47) demonstrates the continuity of the first derivative at $r=R.$
Where the notations $M$ and $\Lambda$ denote the mass and cosmological constant.
It is noteworthy that the cosmological constant $\Lambda$ is contingent upon the constants $\delta_{1}$ and $\delta_{2}$, which are expressed as follows $\Lambda=\frac{\delta_{2}}{2\delta_{1}}$. In this scenario, the evidence for the cosmological constant is zero, hence its value does not affect the current stellar model. Normally, the cosmological constant would have a significant value to influence the mathematical formulation of the problem. The cosmological constant, which is zero in the current study, is estimated to be $10^{-46}km^{-2}$, based on numerical evidence. For this project, $\delta_{2}$ has been set to 0.00001 for the entire problem. $\delta_{1}$ has a significant impact on this proposed model because as we increase the value of $\delta_{1}$, the mathematical values and graphical behaviours of physical properties like density, pressure, adiabatic index, mass, compactness, and velocity of sound also increase. Along with this, the mathematical values of central density and central pressure of the compact star also increase.
Eqs. (4143) are boundary conditions with the help of these boundary conditions we calculated the values of the constants $A_{1}$ and $B_{1}$ that is
@@
$B_{1}=-\frac{\phi_{13}e^{-fR^{2}-h}\left(e^{2\left(fR^{2}+h\right)}-1\right)}{8b\delta_{1}\log(e)}.$
@@ III.3 Behaviour of density
For a well-behaved model of an anisotropic compact star, the density value is maximum at the centre and decreases as the radius increases towards the surface. In our model, the density value is Optimum at the centre and decreases as the radius increases towards the surface as shown in Fig. 2. Therefore our model is well-behaved.
@@ III.5 Anisotropy factor and anisotropic force
In a static and spherically symmetric non-perfect fluid structure, the anisotropy factor, $\Delta=(p_{t}-p_{r})$, defines the dissimilarity in pressures, particularly the radial pressure ($p_{r}$) and the tangential pressure ($p_{t}$).
A force due to the anisotropic pressure is represented by $\Delta/r$, which is repulsive, if $p_{t}>p_{r}\leftrightarrow\Delta>0$, and attractive if $p_{t}<p_{r}\leftrightarrow\Delta<0$ of the stellar model. For the considered matter distribution when $p_{t}>p_{r}$ allows the construction of more compact objects, compared to isotropic fluid sphere Gokhroo:1994fbj ; Maurya:2017uyk ; maurya2017anisotropic ; Maurya:2017jdo ; morales2018compact etc. A positive anisotropy factor improves the stability and equilibrium of the system. The value of $\Delta$ must be zero in the centre of the star structure. Figure 2 indicates that the value of the anisotropy factor $\Delta$ is zero at the centre of the compact star and the value of the anisotropy factor $\Delta$ is positively rising within the star interior of our model, suggesting that the system undergoes a repulsive anisotropic force.
Figure 2: The behaviour of energy density($\rho$), radial pressure($p_{r}$), tangential pressure($p_{t}$) and anisotropy($\Delta$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
Figure 3: The behaviour of gradient of energy density($\frac{d\rho}{dr}$), gradient of radial pressure($\frac{dp_{r}}{dr}$) and gradient of tangential pressure($\frac{dp_{t}}{dr}$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
@@ III.6 Zel’dovich requirement
Figure 4: The behaviour $\omega_{r}$ and $\omega_{t}$ for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
@@ III.7 Energy bounds
When analysing data, all energy constraints must be met for the model to be feasible. Here NEC represents the null energy condition, WEC represents the weak energy condition, DEC represents the dominant energy condition, SEC represents the strong energy condition and TEC depicts the trace energy condition.
Figure 5: The behaviour of NEC, WEC, DEC, SEC and TEC for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
@@
Figure 5 represents the energy requirements graphically; all energy requirements needs have been identified and are fully met. Therefore, our solutions are physically possible.
@@ III.8 Adiabatic index
The adiabatic index is an essential parameter that defines the stability of a stellar structure. For a spherically symmetric object, testing the adiabatic index is crucial since it indicates its solidity. For a stable compact star structure the value of the adiabatic index( $\Gamma$) should be greater than $\frac{4}{3}$ to guarantee stability heintzmann1975neutron .The adiabatic index is often defined as:
@@
Figure 6: Graphical representation of Adiabatic index $\Gamma$ for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
@@ III.9.1 Equilibrium requirement
Figure 7: The behaviour of gravitational force ($F_{g}$), hydrostatic force ($F_{h}$) and anisotropic force ($F_{a}$) versus the radius r of the proposed compact star and $\delta_{1}$ for a fixed value of $\delta_{2}=0.00001$.
@@ III.9.2 Causality and Stability requirement
Figure 8: The behaviour of radial velocity ($V^{2}_{r}$), Tangential velocity ($V^{2}_{t}$ ) and Resultant velocity($|V^{2}_{t}-V^{2}_{r}|$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
@@ III.9.3 Stability Versus Convection
Figure 9: The behaviour of $\rho^{\prime\prime}$ for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
@@ III.9.4 Harrison-Zeldovich-Novikov Stability Condition
Harrison  harrison1965gravitation and Zeldovich-Novikov  zeldovich1971relativistic proposed stability criteria for the compact star model, contingent upon the star’s mass and centre density. According to these observations, they suggested that for a stable configuration, $\frac{dM}{d\rho_{c}}>0$, where M and $\rho_{c}$ represent the mass and central density of the compact star, respectively. With the central density, we will define the mass function of our static anisotropic fluid solution as
@@
Figure 10 demonstrates that our anisotropic models fulfil the Harrison-Zeldovich-Novikov stability criterion. The stability of our structures increases with increasing radii and remains constant after the central densities attain their maximum values.
@@ III.10 Redshift Profile
A phenomenon known as gravitational redshift occurs when electromagnetic waves or photons, lose energy when they emerge from a gravitational well. Electromagnetic radiation expands its wavelength and photon energy to escape when it moves out of a gravitational well. It dissipates energy not by altering its velocity but by varying its frequency, as it must propagate concurrently at the speed of light. When energy is diminished, the photon’s wave frequency declines and its wavelength elongates, leading to a shift toward the red end of the electromagnetic spectrum, referred to as redshift. The behaviour of interior redshift similarly reflects the profile of interior density. A photon must traverse a longer distance and pass through a denser part of the compact star’s interior to exit the centre and reach the surface. This greater dispersion causes a significant energy loss. A photon that ascends near the surface will traverse a shorter distance through a less dense region, leading to less dispersion and energy loss. This means that the interior redshift is highest in the centre and lowest toward the surface. The surface redshift depends on the surface gravity, which is defined by the stellar object’s radius and total mass. Surface gravity increases as a result of a modest increase in radius, similar to the increase in mass. As a result, the surface redshift increases monotonically towards the surface of the star. Gravitational and surface redshifts are defined as:
@@
Where the value of $u(r)$ is given in Eq. (68), the graph of gravitational redshift is monotonically decreasing towards the surface and the graph of surface redshift is monotonically increasing towards the surface as shown in Fig. 11.
Figure 11: The behaviour of Gravitational Redshift ($Z_{G}$) and Surface Redshift ($Z_{S}$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
@@ III.11 Mass and compactness of the model
The compact star’s compactness and mass are shown in Fig. 12. The impact of parameter $\delta_{1}$ on the compactness and mass of the compact star is visible. When the value of $\delta_{1}$ increases, mass and compactness also increase. The mass and compactness values are zero in the centre of the compact star and rise monotonically toward the surface. It is significant to note that there is a mass-radius ratio bound for any stellar structure, which is provided by Buchdahl and is $\frac{2M}{R}\leqslant\frac{8}{9}$ Buchdahl:1959zz This bound has been computed for the compact star in our model and is displayed in Table 1. This bound is less than $\frac{4}{9}$, demonstrating that our model satisfies Buchdahl’s requirements. The functions $m(r)$, $u(r)$ of radial coordinate r are used to represent the compact star mass and compactness are given below:
@@
The star obtained here is inside the star’s limit of being a compact star, with a solar mass of 1.49 for a value of $\delta_{1}$=1. The graph’s growing tendency suggests that the star’s mass rises gradually as the parameter $\delta_{1}$ rises. In this scenario, the constant $\delta_{2}$ is supposed to be 0.00001, therefore it has minimal impact on the mass and compactness of the model.
Figure 12: The behaviour of mass function $m(r)$ and compactness parameter $u(r)$ versus the radius r of the proposed compact star and $\delta_{1}$ for a fixed value of $\delta_{2}=0.00001$.

### v2

@@ II.1 Field equations of $\mathcal{F}(\mathcal{Q})$ gravity
where, $\delta_{1}$ and $\delta_{2}$ are integration constants. Using equations Eq. (25) and Eq. (26), the field equations for $\mathcal{F}(\mathcal{Q})$ gravity can be written as:
@@
$p_{r}=\frac{1}{16\pi r^{2}}\Bigg{[}-2\delta_{1}+2e^{-\sigma}\delta_{1}(r\Upsilon^{\prime}+1)+r^{2}\delta_{2}\Bigg{]},$
$p_{t}=\frac{e^{-\sigma}}{32\pi r}\Bigg{[}2e^{\sigma}r\delta_{2}+\delta_{1}(2+r\Upsilon^{\prime})(\Upsilon^{\prime}-\sigma^{\prime})+2r\delta_{1}\Upsilon^{\prime\prime}\Bigg{]}.$
In Equations (2929) The prime symbol ()($\prime$)( ′ ) signifies differentiation with respect to the radial coordinate $r$, whereas $\rho$, $p_{r}$, $p_{t}$, $\sigma$, and $\Upsilon$ are variables that we will resolve to get the desired outcome. For static and SS fluid systems, the generic formulation for the anisotropic component, written as $\Delta=(p_{t}-p_{r})$, is articulated as follows:
@@ II.2 Karmarkar Condition
The SS line element (18) may continually be embedded in six-dimensional flat spacetime, signifying its classification as belonging to embedding class two in general. On the contrary, the SS metric can be included in five-dimensional flat spacetime, provided it fulfils the Karmarkar criterion  karmarkar1948gravitational . Then It indicates the spacetime of embedding class one. It is a necessary and sufficient condition for SS spacetime to belong to class one. The Karmarkar condition is expressed through the curvature components as follows:
@@ III.1 Metric elements
Anisotropic compact stars require finite and nonsingular metric elements to produce a physically plausible model. According to our solution, $e^{\sigma}$ at $r=0$ equals 1 and $e^{\Upsilon}$ at r=0 equals 0.4583. At r=0, we see a positive value of $e^{\sigma}$ and $e^{\Upsilon}$, showing that the metric components are consistent at the core of the object. The precise behavior of the metric components is displayed in Fig. 1. The graphs of the components $e^{-\sigma}$ and $e^{\Upsilon}$ are bounded and nonsingular within the stellar configuration. The graph of $e^{-\sigma}$ exhibits a monotonically decreasing trend near the surface, whereas the graph of $e^{\Upsilon}$ demonstrates a monotonically increasing trend approaching the surface. The graph intersects near the border of the compact star; given these qualities of the metric components, we may infer that they are capable of generating a model for the anisotropic stellar model.
@@ III.2 Seamless Matching of Interior and Exterior spacetime
This section presents the boundary condition necessary for resolving the Einstein field equations in $\mathcal{F}(\mathcal{Q})$ gravity. To do this, we compare the solutions of the outer and inner spacetimes. Recent studies, such as Wang:2021zaz , suggest that the viable solution in $\mathcal{F}(\mathcal{Q})$ gravity may be resolved using the Schwarzschild (Anti-) de Sitter solution. The result is presented as follows:
@@
By setting $r=R$, the Schwarzschild (Anti-)de Sitter metric has been solved at the boundary in accordance with the criteria established by Darmois-Israel in their study  Israel:1966rt as:
@@
The analysis from Eq. (46) and Eq. (47) demonstrates the continuity of the first derivative at $r=R.$ The symbols $M$ and $\Lambda$ represent the mass and the cosmological constant, respectively.
The cosmological constant $\Lambda$ is dependent on the constants $\delta_{1}$ and $\delta_{2}$, represented by the equation $\Lambda=\frac{\delta_{2}}{2\delta_{1}}$. In this context, the evidence for the cosmological constant is nonexistent; hence, its value does not influence the present star model. The cosmological constant often has a substantial value that affects the mathematical formulation of the issue. The cosmological constant, currently assessed as zero in this research, is projected to be $10^{-46}km^{-2}$ according to numerical evidence. In this project, $\delta_{2}$ has been established as 0.00001 throughout the duration of the problem. The parameter $\delta_{1}$ significantly influences the suggested model; an increase in $\delta_{1}$ results in corresponding increases in the mathematical values and graphical behaviors of physical parameters such as density, pressure, adiabatic index, mass, compactness, and sound velocity. Simultaneously, the center density and central pressure of the stellar model also increase.
Equations (4143) represent boundary conditions, which facilitated the calculation of the constants $A_{1}$ and $B_{1}$ as,
@@
$B_{1}=-\frac{\phi_{13}e^{-fR^{2}-h}\left(e^{2\left(fR^{2}+h\right)}-1\right)}{8b\delta_{1}\log(e)}.$ (49)
@@ III.3 Behaviour of density
For a well-behaved model of an anisotropic compact star, the density value is maximum at the centre and decreases as the radius increases towards the surface. In our model, the density value is optimum at the centre and decreases as $r$ rises as shown in Fig. 2. Therefore our model is well-behaved.
@@ III.5 Anisotropy factor and anisotropic force
In a static and SS non-ideal fluid system, the anisotropy factor, $\Delta=(p_{t}-p_{r})$, establishes the inequality in pressures, namely $p_{r}$ and $p_{t}$. A force resulting from anisotropic pressure is denoted by $\Delta/r$, which becomes repulsive when $p_{t}>p_{r}\leftrightarrow\Delta>0$, and attractive when $p_{t}<p_{r}\leftrightarrow\Delta<0$ in the star model. In the examined matter distribution, when $p_{t}>p_{r}$, it facilitates the formation of more compact objects in comparison to an isotropic fluid sphere Gokhroo:1994fbj ; Maurya:2017uyk ; maurya2017anisotropic ; Maurya:2017jdo ; morales2018compact . The system’s stability and equilibrium are enhanced by a positive anisotropy factor. In the center of the star, $\Delta$ has to equal zero. The system remains a repulsive anisotropic force, as shown by Figure 2, where the value of the $\Delta$ is positively increasing throughout the star and zero at the center of the star.
Figure 2: The behaviour of energy density($\rho$), radial pressure($p_{r}$), tangential pressure($p_{t}$) and anisotropy($\Delta$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and $r$.
Figure 3: The behaviour of gradient of energy density($\frac{d\rho}{dr}$), gradient of radial pressure($\frac{dp_{r}}{dr}$) and gradient of tangential pressure($\frac{dp_{t}}{dr}$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and $r$.
@@ III.6 Zel’dovich requirement
Figure 4: The behaviour $\omega_{r}$ and $\omega_{t}$ for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and $r$.
@@ III.7 Energy bounds
All energy limitations must be satisfied for the model to be viable during data analysis. NEC denotes the null energy condition, WEC signifies the weak energy condition, DEC indicates the dominant energy condition, SEC denotes the strong energy condition, and TEC illustrates the trace energy condition.
Figure 5: The behaviour of NEC, WEC, DEC, SEC and TEC for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and $r$.
@@
Figure 5 represents the energy requirements graphically; All energy constraints have been determined and are completely satisfied. Consequently, our solutions are physically feasible.
@@ III.8 Adiabatic index
The adiabatic index is an essential parameter that defines the stability of a stellar model. For a SS object, testing the adiabatic index is crucial since it indicates its solidity. For a stable compact star structure the value of the adiabatic index( $\Gamma$) should be greater than $\frac{4}{3}$ to guarantee stability heintzmann1975neutron .The adiabatic index is often defined as:
@@
Figure 6: Graphical representation of Adiabatic index $\Gamma$ for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and $r$.
@@ III.9.1 Equilibrium requirement
Figure 7: The behaviour of gravitational force ($F_{g}$), hydrostatic force ($F_{h}$) and anisotropic force ($F_{a}$) versus $r$ and $\delta_{1}$ for a fixed value of $\delta_{2}=0.00001$.
@@ III.9.2 Causality and Stability requirement
Figure 8: The behaviour of radial velocity ($V^{2}_{r}$), Tangential velocity ($V^{2}_{t}$ ) and Resultant velocity($|V^{2}_{t}-V^{2}_{r}|$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and $r$.
@@ III.9.3 Stability Versus Convection
Figure 9: The behaviour of $\rho^{\prime\prime}$ for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and $r$.
@@ III.9.4 Harrison-Zeldovich-Novikov Stability Condition
Depending on the star’s mass and center density, Harrison  harrison1965gravitation and Zeldovich-Novikov  zeldovich1971relativistic suggested stability requirements for the compact star model. Based on this information, they proposed that $\frac{dM}{d\rho_{c}}>0$ for a stable structure, where M and $\rho_{c}$ stand for the compact star’s mass and central density, respectively. The mass function of our solution can be defined using the center density as
@@
Figure 10 illustrates that our anisotropic models satisfy the Harrison-Zeldovich-Novikov stability criteria. The stability of our constructions improves with larger radii and stays constant after the central densities reach their peak levels.
@@ III.10 Redshift Profile
Gravitational redshift is a phenomena in which electromagnetic waves or photons lose energy as they ascend from a gravitational well. Electromagnetic radiation grows its wavelength and diminishes photon energy when it exits a gravitational well. Since it must propagate simultaneously at the speed of light, it consumes energy by changing its frequency rather than its velocity. When energy is reduced, the photon’s wave frequency decreases and its wavelength extends, resulting in a shift toward the red end of the electromagnetic spectrum, known as redshift. The behaviour of interior redshift similarly reflects the profile of interior density. A photon must traverse a longer distance and pass through a denser part of the compact star’s interior to exit the centre and reach the surface. This greater dispersion causes a significant energy loss. A photon that ascends near the surface will traverse a shorter distance through a less dense region, leading to less dispersion and energy loss. The inner redshift is maximal at the center and minimal at the surface. The surface redshift is contingent upon the surface gravity, which is determined by the radius and total mass of the star object. Surface gravity rises due to a little augmentation in radius, equivalent to the growth in mass. Consequently, the surface redshift escalates monotonically approaching the star’s surface. Gravitational and surface redshifts are characterized as follows:
@@
The value of $u(r)$ is specified in Eq. (68), indicating that the gravitational redshift monotonically decreases approaching the surface, whereas surface redshift monotonically increases when we move to surface, as seen in Fig. 11.
Figure 11: The behaviour of Gravitational Redshift ($Z_{G}$) and Surface Redshift ($Z_{S}$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and $r$.
@@ III.11 Mass and compactness of the model
The mass and compactness of the compact star are shown in Fig. 12. It is possible to see how parameter $\delta_{1}$ affects the compactness and mass of the compact star. Increasing $\delta_{1}$ results in an increase in mass and compactness. The compact star’s mass and compactness values are zero in its core and increase monotonically as it moves toward the surface. It’s important to remember that every star structure has a mass-radius ratio constraint, which Buchdahl provides and is $\frac{2M}{R}\leqslant\frac{8}{9}$. In Ref. Buchdahl:1959zz , this limit is shown in Table 1 and has been calculated for the our compact star model. This constraint is inferior than $\frac{4}{9}$, indicating that our model adheres to Buchdahl’s criteria. The functions $m(r)$ and $u(r)$, corresponding to the radial coordinate $r$, reflect the mass and compactness of the compact star, respectively, as shown below:
@@
The star acquired here lies inside the threshold of being a compact star, with a solar mass of 1.49 for a value of $\delta_{1}$=1. The graph’s upward trend indicates that the star’s mass increases progressively with the rising of the parameter $\delta_{1}$. In this context, the constant $\delta_{2}$ is set to 0.00001, thereby exerting little influence on the mass and compactness of the model.
Figure 12: The behaviour of $m(r)$ and $u(r)$ versus and $r$ and $\delta_{1}$ for a fixed value of $\delta_{2}=0.00001$.